OpenStudy (anonymous):

determine whether (6,-2) is a solution of the system x-y=8 and 2x+3y=6

OpenStudy (anonymous):

plug in x =6 and y = -2 in the system and see if they satisfy the system

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

because both the equation is satisfied with the point6,-2

OpenStudy (anonymous):

i got no

OpenStudy (anonymous):

putting the point in the lhs of the 1st equation 6-(-2)=6+2=8 which is the rhs.....similar for the second equation

OpenStudy (anonymous):

then you substituted incorrectly. try again, and pay attention to addition and subtraction

OpenStudy (anonymous):

oh ok i see wat i did wrong